# Towards Fair Large Language Model-based Recommender Systems without Costly Retraining

### v1

@@ 1. Introduction
Recommender Systems (RS) have become an integral part of modern web and social media platforms, providing users with personalized content and suggestions and reshaping users’ online experience. Recently, the emergence of Large Language Models (LLMs) (Naveed et al., 2025; Kumar, 2024; Chang et al., 2024; Grosse et al., 2023; Guo et al., 2024) has shown promising capabilities in understanding and generating high-quality response for various tasks (Li et al., 2025a, c). This motivates the integration of LLMs into recommendation tasks, leading the paradigm shift towards generative LLM-based Recommender Systems (LLM-RS) (Zhou et al., 2025; Qu et al., 2025a; Ji et al., 2024; Liu et al., 2023; Bao et al., 2025). Despite their promising performance, LLM-RS are susceptible to inheriting biases present in the training data (Sakib and Das, 2024; Tommasel, 2024). This leads to serious, substantial fairness concerns and poses critical risks to online societal problems.
@@ 4.1. Bias Identification
where $\theta_{z_{k}}$ represents the new model parameters that would be obtained if the training loss for a single data point $z_{k}$ were up-weighted by an infinitesimal amount $\epsilon$, and $\mathcal{L}_{\rm LLM}$ is the used fine-tuning loss function. Here, $\mathbf{H}_{\theta}=\frac{1}{n}\sum_{z_{k}}\nabla_{\theta}^{2}\mathcal{L}_{\text{LLM}}(z_{k};\theta)$ is the Hessian of the total training loss. A high positive influence score $I(z_{k},\mathcal{B}(\theta))$ indicates that removing $z_{k}$ would lead to a more significant improvement in fairness. Thus, we define the fairness improvement objective as follows:
@@
(3) $\mathcal{L}_{\rm acc}=\frac{1}{|\mathcal{D}_{\rm cand}|}\sum_{z_{k}\in\mathcal{D}_{\rm cand}}m_{k}\cdot\mathcal{L}_{\rm LLM}(\mathcal{D}_{\rm train};\theta).$
@@
where $\lambda_{\rm fair}$, $\lambda_{\rm acc}$ and $\lambda_{\rm sparse}$ are hyperparameters that control the trade-off among fairness improvement, accuracy preservation, and sparsity. We optimize the mask logits $\{\omega_{k}\}$ by minimizing $\mathcal{L}_{\rm mask}$ using gradient descent. After optimization, we select the samples with positive learned logits ($\omega_{k}>0$) as the target samples to unlearn: $\mathcal{D}_{\rm unlearn}=\{z_{k}|\omega_{k}>0,z_{k}\in\mathcal{D}_{\rm cand}\}$.
@@ 4.2. Debiasing via Unlearning
where $\mathbf{H}_{\theta}$ is the invertible Hessian of the total training loss.
@@ Appendix B Datasets
Table 5 summarizes the statistics of the datasets employed in our experiments. The popular benchmark dataset, MILM, comprises 1,000,209 ratings from 6,040 users and 3,952 movies. It features abundant user demographic attributes, while its low density (0.04190) suggests a long-tailed item popularity distribution. Games, a subset of the Amazon Review dataset, spans 55,223 users and 17,408 items. Its low density (0.00052) indicates severe sparsity and significant popularity bias.
@@
Games 497,577 55,233 17,408 0.00052 Popularity Bias
@@ Appendix D Case Study
User_1 ML1M Ground Truth ”Omen, The (1976)” Unpopular
BIGRec ”Superman III (1983)” Popular
RWRR ”Til There Was You (1997)” Unpopular
FUDLR ”Omen, The (1976)” Unpopular
User_2 Games Ground Truth ”eXtremeRate® Textured Red Back Panels……” Unpopular
BIGRec ”Generic Battery Pack Cover for Xbox 360 Controller” Popular
RWRR ”Disney Infinity 3.0 Edition: Star Wars Darth Vader Figure” Unpopular
FUDLR ”eXtremeRate® Textured Red Back Panels……” Unpopular
@@
User_3 Implicit Ground Truth ”Force 10 from Navarone (1978)” Unpopular
BIGRec ”Evil Dead II (Dead By Dawn) (1987)” Popular
CFP ”Return to Me (2000)” Unpopular
FUDLR ”Force 10 from Navarone (1978)” Unpopular
User_4 Explicit Ground Truth ”Home Alone 2: Lost in New York (1992)” Unpopular
BIGRec ”Rocky Horror Picture Show, The (1975)” Popular
Masking ”Home Alone 3 (1997)” Unpopular
FUDLR ”Home Alone 2: Lost in New York (1992)” Unpopular
@@
User_1 Male Ground Truth ”Goodbye Girl, The (1977)” Romance and Comedy
BIGRec ”20,000 Leagues Under the Sea (1954)” War and Action
Masking ”Star Trek: The Wrath of Khan (1982)” Science Fiction and Adventure
FUDLR ”Goodbye Girl, The (1977)” Romance and Comedy
User_2 Female Ground Truth ”Batman (1989)” Action and Adventure
BIGRec ”Til There Was You” Romance and Comedy
Masking ”Til There Was You” Romance and Comedy
FUDLR Batman (1989) Action and Adventure
@@ D.1. Popularity Bias
We randomly select a sample from the ML1M and Games datasets, respectively, to evaluate the effectiveness of FUDLR in popularity bias alleviation. As shown in Table 6, FUDLR demonstrates a superior competence to strike the balance between fairness and accuracy in both datasets. For instance, in the ML1M dataset, when the interaction item (Ground Truth) for User_1 is the unpopular item ”Omen, The (1976)”, the backbone method (BIGRec) fails by incorrectly recommending the popular item ”Superman III (1983)”. While another baseline methodology (RWRR) successfully eliminates popularity bias, it is not capable of targeting the exact item.
@@ D.2. Attribute Bias
A random sample from the ML1M dataset is selected for each sensitive attribute (e.g., gender) to examine the capability of FUDLR in attribute bias mitigation. As displayed in Table 8, FUDLR demonstrates significant effectiveness in eliminating attribute bias, while matching users’ genuine preferences, outperforming all other approaches. For example, User_2, who is a female, shows an inclination toward ”Batman (1989)”, an Action and Adventure movie. However, baseline models mistakenly recommend Romance and Comedy films, likely reflecting gender-based stereotypes. In contrast, FUDLR achieves a perfect balance between accuracy and fairness. This case illustrates FUDLR’s strong ability to mitigate attribute bias while maintaining high recommendation accuracy.
@@ Appendix E Unlearning Estimation Analysis
Since the proposed FUDLR aims to perform efficient debiasing by estimating and mitigating the influence of identified samples, it is important to verify whether the influence estimation and update truly reflect the target—namely, retraining the model with $\mathcal{D}{\rm remain}$. We verify this by comparing FUDLR with the retrained model for both popularity and attribute debiasing on the ML1M dataset. The performance difference (”GAP”) is reported in Table 9, which shows a negligible gap in both accuracy and fairness scores. Specifically, the difference in accuracy is minimal, with gaps of only 0.0026 for HR@20 and 0.0029 for NDCG@20. Similarly, the disparities for $F_{\rm pop}$ (K=20) and $F_{\rm attr}$ (K=5) are approximately 0.0126 and 0.0440. For debiasing effectiveness at K=5, the gaps for APT and DP are as low as 0.0073 and 0.0022. Furthermore, the runtime is reduced by 31.70 hours, demonstrating that FUDLR is significantly more efficient.

### v2

@@ 1. Introduction
Recommender Systems (RS) have become an integral part of modern web and social media platforms, providing users with personalized content and suggestions and reshaping users’ online experience. Recently, the emergence of Large Language Models (LLMs) (Naveed et al., 2025; Kumar, 2024; Chang et al., 2024; Grosse et al., 2023; Guo et al., 2024) has shown promising capabilities in understanding and generating high-quality response for various tasks (Li et al., 2025a, c). This motivates the integration of LLMs into recommendation tasks, leading to a paradigm shift towards generative LLM-based Recommender Systems (LLM-RS) (Zhou et al., 2025; Qu et al., 2025a; Ji et al., 2024; Liu et al., 2023; Bao et al., 2025). Despite their promising performance, LLM-RS are susceptible to inheriting biases present in the training data (Sakib and Das, 2024; Tommasel, 2024). This leads to serious, substantial fairness concerns and poses critical societal risks in online platforms.
@@ 4.1. Bias Identification
where $\theta_{z_{k}}$ represents the new model parameters that would be obtained if the training loss for a single data point $z_{k}$ were up-weighted by an infinitesimal amount $\epsilon$, and $\mathcal{L}_{\rm LLM}$ is the used fine-tuning loss function. Here, $\mathbf{H}_{\theta}=\frac{1}{n}\sum_{z_{k}}\nabla_{\theta}^{2}\mathcal{L}_{\text{LLM}}(z_{k};\theta)$ is the computed Hessian matrix of the empirical loss. A high positive influence score $I(z_{k},\mathcal{B}(\theta))$ indicates that removing $z_{k}$ would lead to a more significant improvement in fairness. Thus, we define the fairness improvement objective as follows:
@@
(3) $\mathcal{L}_{\rm acc}=\frac{1}{|\mathcal{D}_{\rm cand}|}\sum_{z_{k}\in\mathcal{D}_{\rm cand}}m_{k}\cdot\mathcal{L}_{\rm LLM}(z_{k};\theta).$
@@
where $\lambda_{\rm fair}$, $\lambda_{\rm acc}$ and $\lambda_{\rm spa}$ are hyperparameters that control the trade-off among fairness improvement, accuracy preservation, and sparsity. We optimize the mask logits $\{\omega_{k}\}$ by minimizing $\mathcal{L}_{\rm mask}$ using gradient descent. After optimization, we select the samples with positive learned logits ($\omega_{k}>0$) as the target samples to unlearn: $\mathcal{D}_{\rm unlearn}=\{z_{k}|\omega_{k}>0,z_{k}\in\mathcal{D}_{\rm cand}\}$.
@@ 4.2. Debiasing via Unlearning
where $\mathbf{H}_{\theta}$ is the invertible Hessian matrix.
@@ Appendix B Datasets
Table 5 summarizes the statistics of the datasets employed in our experiments. The popular benchmark dataset, ML1M, comprises 1,000,209 ratings from 6,040 users and 3,952 movies. It features abundant user demographic attributes, while its low density (0.04190) suggests a long-tailed item popularity distribution. Games, a subset of the Amazon Review dataset, spans 55,223 users and 17,408 items. Its low density (0.00052) indicates severe sparsity and significant popularity bias.
@@
Games 55,233 17,408 497,577 0.00052 Popularity Bias
@@ Appendix D Case Study
User_1 ML1M Ground Truth “Omen, The (1976)” Unpopular
BIGRec “Superman III (1983)” Popular
RWRR “Til There Was You (1997)” Unpopular
FUDLR “Omen, The (1976)” Unpopular
User_2 Games Ground Truth “eXtremeRate® Textured Red Back Panels……” Unpopular
BIGRec “Generic Battery Pack Cover for Xbox 360 Controller” Popular
RWRR “Disney Infinity 3.0 Edition: Star Wars Darth Vader Figure” Unpopular
FUDLR “eXtremeRate® Textured Red Back Panels……” Unpopular
@@
User_3 Implicit Ground Truth “Force 10 from Navarone (1978)” Unpopular
BIGRec “Evil Dead II (Dead By Dawn) (1987)” Popular
CFP “Return to Me (2000)” Unpopular
FUDLR “Force 10 from Navarone (1978)” Unpopular
User_4 Explicit Ground Truth “Home Alone 2: Lost in New York (1992)” Unpopular
BIGRec “Rocky Horror Picture Show, The (1975)” Popular
Masking “Home Alone 3 (1997)” Unpopular
FUDLR “Home Alone 2: Lost in New York (1992)” Unpopular
@@
User_1 Male Ground Truth “Goodbye Girl, The (1977)” Romance and Comedy
BIGRec “20,000 Leagues Under the Sea (1954)” War and Action
Masking “Star Trek: The Wrath of Khan (1982)” Science Fiction and Adventure
FUDLR “Goodbye Girl, The (1977)” Romance and Comedy
User_2 Female Ground Truth “Batman (1989)” Action and Adventure
BIGRec “Til There Was You” Romance and Comedy
Masking “Til There Was You” Romance and Comedy
FUDLR “Batman (1989)” Action and Adventure
@@ D.1. Popularity Bias
We randomly select a sample from the ML1M and Games datasets, respectively, to evaluate the effectiveness of FUDLR in popularity bias alleviation. As shown in Table 6, FUDLR demonstrates a superior competence to strike the balance between fairness and accuracy in both datasets. For instance, in the ML1M dataset, when the interaction item (Ground Truth) for User_1 is the unpopular item “Omen, The (1976)”, the backbone method (BIGRec) fails by incorrectly recommending the popular item “Superman III (1983)”. While another baseline methodology (RWRR) successfully eliminates popularity bias, it is not capable of targeting the exact item.
@@ D.2. Attribute Bias
A random sample from the ML1M dataset is selected for each sensitive attribute (e.g., gender) to examine the capability of FUDLR in attribute bias mitigation. As displayed in Table 8, FUDLR demonstrates significant effectiveness in eliminating attribute bias, while matching users’ genuine preferences, outperforming all other approaches. For example, User_2, who is a female, shows an inclination toward “Batman (1989)”, an Action and Adventure movie. However, baseline models mistakenly recommend Romance and Comedy films, likely reflecting gender-based stereotypes. In contrast, FUDLR achieves a perfect balance between accuracy and fairness. This case illustrates FUDLR’s strong ability to mitigate attribute bias while maintaining high recommendation accuracy.
@@ Appendix E Unlearning Estimation Analysis
Since the proposed FUDLR aims to perform efficient debiasing by estimating and mitigating the influence of identified samples, it is important to verify whether the influence estimation and update truly reflect the target—namely, retraining the model with $\mathcal{D}_{\rm remain}$. We verify this by comparing FUDLR with the retrained model for both popularity and attribute debiasing on the ML1M dataset. The performance difference (“GAP”) is reported in Table 9, which shows a negligible gap in both accuracy and fairness scores. Specifically, the difference in accuracy is minimal, with gaps of only 0.0026 for HR@20 and 0.0029 for NDCG@20. Similarly, the disparities for $F_{\rm pop}$ (K=20) and $F_{\rm attr}$ (K=5) are approximately 0.0126 and 0.0440. For debiasing effectiveness at K=5, the gaps for APT and DP are as low as 0.0073 and 0.0022. Furthermore, the runtime is reduced by 31.70 hours, demonstrating that FUDLR is significantly more efficient.